9% of population has a symptom of a disease. If one tests a symptomatic person, probability of testing positive is p=0.78. If one tests a person without symptoms, the probability is p=0.06. If you test a random person what is the probability of the test returning positive results? What is the probability of a positive testing person actually having a disease?
Answer by ikleyn(52780)

.
9% of population has a symptom of a disease.
If one tests a symptomatic person, probability of testing positive is p=0.78.
If one tests a person without symptoms, the probability is p=0.06.
(a) If you test a random person what is the probability of the test returning positive results?
(b) What is the probability of a positive testing person actually having a disease?
~~~~~~~~~~~~~~~

(a)  A random person EITHER has symptoms of the disease with the probability 0.09 
                       OR   has no symptoms with the probability of 1 - 0.09 = 0.91.


     If the person has symptoms of the desease,    then he contributes 0.09*0.78 to the probability to get a positive test.

     If the person has no symptoms of the desease, then he contributes 0.91*0.06 to the probability to get a positive test.


     So, the total probability to get the positive test is 

         P = 0.09*0.78 + 0.91*0.06 = 0.1248.    ANSWER



(b)  This probability is 0.78, as stated in the problem.    ANSWER

Solved.